OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (−4, 8) and (−3, −6). Choose one answer. a. 7 b. −14 c. zero d. 14

OpenStudy (anonymous):

must find slope first. equation for slope is: y2-y1/x2-x1 plug in the points: -6-8/-3- -4 = -14/1 = -14 so slope = -14

OpenStudy (anonymous):

so the answer is b.